OpenStudy (anonymous):

This limit is driving me crazy Lim x^2COSX x->0 --------- 2Sin^2(x/2)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ x ^{2}cosx}{ 2\sin^2(x/2) } \]

OpenStudy (anonymous):

does your question is like this????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what always gets to me are the COS and SIn

OpenStudy (anonymous):

Im supposed to calculate it but im having trouble with the COS and the SIN

OpenStudy (anonymous):

ok wait let me solve it then i will explain it to you

OpenStudy (anonymous):

if you substitute 0 for x you will get 0/0 form which is undefined case right

OpenStudy (anonymous):

yes i understand

OpenStudy (anonymous):

so you have to look for the denominator which is 2(sin(x/2))^2 now use half angle trigonometric identity

OpenStudy (anonymous):

[sin(x/2)]^2 = 1/2[1-cosx]

OpenStudy (anonymous):

putting this denomenator you will get

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ x^2cosx }{ 2.1/2.(1-cosx) } \]

OpenStudy (anonymous):

2 . (1/2)(1-COSX) right?

OpenStudy (anonymous):

ya now cancel 2 with 1/2

OpenStudy (anonymous):

And it becomes one leaving me with x^2COSX --------- (1-COSX)

OpenStudy (anonymous):

at this point i m stuck a little plz wait

OpenStudy (turingtest):

use\[\lim_{x\to0}{\sin x\over x}=1\]

OpenStudy (turingtest):

\[\lim_{x\to0}{x^2\cos x\over\sin^2(x/2)}=\lim_{x\to0}\cos x\cdot\lim_{x\to0}\left({1\over\frac{\sin (x/2)}x}\right)^2\]\[=4\lim_{x\to0}\cos x\cdot\lim_{x\to0}\left({1\over\frac{\sin (x/2)}{x/2}}\right)^2\]

OpenStudy (turingtest):

@Chef86 does this make sense to you?

OpenStudy (anonymous):

@TuringTest you missed a 2 in denominator in frist step, rest is fine

OpenStudy (anonymous):

true the 2 is missing

OpenStudy (turingtest):

ah you are right :)

OpenStudy (turingtest):

\[\lim_{x\to0}{x^2\cos x\over2\sin^2(x/2)}=2\lim_{x\to0}\cos x\cdot\lim_{x\to0}\left({1\over\frac{\sin (x/2)}{x/2}}\right)^2\]

OpenStudy (turingtest):

is there part of that you are having trouble with?

OpenStudy (anonymous):

The Sin(x/2) after dividing would be Sin1 right

OpenStudy (anonymous):

@TuringTest thanks for the info

OpenStudy (turingtest):

@Chef86 I don't think I understand what you mean by "after dividing" @03453660 welcome, please help Chef understand if you could, I have to leave Bye y'all

OpenStudy (anonymous):

leme solve it for u

OpenStudy (anonymous):

from where to start?

OpenStudy (anonymous):

Where Turing left off.

OpenStudy (anonymous):

now apply limits to get result

OpenStudy (anonymous):

is it fine now?